I am stuck on 3 math problems I have left for homework. COuld you please show me how to figure them out? thank you
I am supposed to define a variable, write the equation and solve.
Question:
The difference of 2 numbers is 12.
2/5th of the greatest number is 6 more than 1/3 of the lesser number. Find both numbers.

Answer by stanbon(75887)

I am supposed to define a variable, write the equation and solve.
Question:
The difference of 2 numbers is 12.
2/5th of the greatest number is 6 more than 1/3 of the lesser number. Find both numbers.
---------
Let one of the numbers be "x".
The other number is "x+12"
EQUATION:
(2/5)(x+12)=(1/3)x + 6
Multiply thru by 15 to get:
6(x+12)=5x+90
6x+72=5x+90
x=18 (the smaller number)
x+12=30 (the larger number)
Cheers,
Stan H.